HUBBLE SPACE TELESCOPE DAILY REPORT #1243 PERIOD COVERED: 0000Z Daily Status Report as of 286/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 5498 (Flickering as a Probe of Accretion Disk Structures) The Faint Object Spectrograph (Blue Side) was used in the rapid mode to obtain spectra of the cataclysmic variable star V603-AQL. "Flickering" has been observed in stars of this type due to the accretion process. The flickering rate will be compared in the ultraviolet and longer wavelengths. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 5371 (HST Medium_ Deep Survey: Cycle 4 Medium Low Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make low galactic latitude deep multicolor images of objects such as galaxies, and stars in our galaxy. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of ESOP138-G1 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Four Sets of FOS/BL 5444 (A UV Spectral Library of Metal- Poor Massive Stars) The FOS (Blue Side) was used for a snapshot program to obtain ultraviolet spectra of BI170, a metal-poor star in the LMC. It will be useful for stellar atmosphere and stellar evolution studies. The observations completed without incident. 1.5 Completed FOS/BL 5676 (UV Spectroscopy and Imaging in the Crab Nebula) The Faint Object Spectrograph (Blue Side) was used to obtain spectra of several knots and filaments (WM09-0542) in the Crab supernova remnant to refine the estimates the chemical abundances which are needed for comparison with stellar evolution and the nucleosynthesis theory. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Significant HSTARs: None 2.4 Operations Notes: HST operations personnel report that there were no anomalies in any onboard susbsystem. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-